# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

### v2

@@ B.1 Bessel function identities
In this subsection, we list specific identities of cylindrical Bessel functions $J_{\nu}(z)$ of order $\nu$ which are used throughout our work [25].
$J_{\nu}(z)$ $=\frac{2(\nu+1)}{z}J_{\nu+1}(z)-J_{\nu+2}(z)$ (2.1)
$J_{\nu}(z)$ $=\frac{\big{(}4(\nu+1)(\nu+2)-z^{2}\big{)}J_{\nu+2}(z)-2z(\nu+1)J_{\nu+3}(z)}{z^{2}}$ (2.2)
$\frac{\partial(z^{-\nu}J_{\nu}(z))}{\partial z}$ $=-z^{-\nu}J_{\nu+1}(z)$ (2.3)
@@ B.2 Inverse Laplace transform identities
where $\xi$ is some real constant which, in this paper, only ever takes values $\xi=1,0,-1,-2$, while $\beta$ [not to be confused with the inverse of the temperature $(k_{B}T)^{-1}$] and the chemical potential $\mu$ are conjugate variables of the Laplace transformation. Going between the first and second lines in the derivation of equation 2.4, we utilized three common inverse Laplace transforms which are listed below [10, 25], and going between the third and fourth lines we made the substitution $\mu-V(\mathbf{R})=\frac{\hbar^{2}}{2m}k_{F}(\mathbf{R})^{2}$, suppressed the $\mathbf{R}$ dependent notation of $k_{F}$ so that $k_{F}=k_{F}(\mathbf{R})$, and defined $z=k_{F}s$.
@@ B.3 Gradient identities
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ $=\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}.$ (2.10)
As is the case throughout the rest of the paper, an equal ($=$) sign denotes an exact identity, and an approximation sign ($\approx$) denotes an approximate one, where only terms up to and including second order in $\mathbf{s}$ (equivalently $\nabla_{\mathbf{R}}$ and $\hbar$) are retained. Equations 2.5, 2.6 and 2.7 are straightforward to show via equation 1.3 and the definition of $z$, but equations 2.8, 2.9 and 2.10 are more complicated and deserve their own proofs. Starting with equation 2.8 we show that
$\nabla_{\mathbf{r}}V(\mathbf{r})=\nabla_{\mathbf{r}}V(\mathbf{R}+\mathbf{s}/2)$ $\approx\nabla_{\mathbf{r}}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{r}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$\approx\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{s}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$=\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$=\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)},$
where equation 2.12 was applied in the first line and equation 2.5 between lines one and two. For equation 2.9, we find
@@
$\approx\nabla_{\mathbf{r}}^{2}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{r}}^{2}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{r}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$\approx\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{8}\nabla_{\mathbf{s}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}+\frac{1}{2}\big{(}\nabla_{\mathbf{R}}\cdot\nabla_{\mathbf{s}}\big{)}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$=\frac{3}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{s}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$=\frac{3}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})$
$=\nabla_{\mathbf{R}}^{2}V(\mathbf{R}),$
where equation 2.12 was applied in the first line and equation 2.6 in between lines one and two. For equation 2.10, it is useful write gradients and their dot products in terms of sums and products of vector components. We find the following

### v3

@@ B.1 Bessel function identities
In this subsection, we list specific identities of cylindrical Bessel functions $J_{\nu}(z)$ of order $\nu$ which are used throughout our work [28],
$J_{\nu}(z)$ $=\frac{2(\nu+1)}{z}J_{\nu+1}(z)-J_{\nu+2}(z),$ (2.1)
$J_{\nu}(z)$ $=\frac{\big{(}4(\nu+1)(\nu+2)-z^{2}\big{)}J_{\nu+2}(z)-2z(\nu+1)J_{\nu+3}(z)}{z^{2}},$ (2.2)
$\frac{\partial(z^{-\nu}J_{\nu}(z))}{\partial z}$ $=-z^{-\nu}J_{\nu+1}(z).$ (2.3)
@@ B.2 Inverse Laplace transform identities
where $\xi$ is some real constant which, in this paper, only ever takes values $\xi=1,0,-1,-2$, while $\beta$ (not to be confused with the inverse of the temperature) and the Fermi energy $\mu$ are conjugate variables of the Laplace transformation. Going between the first and second lines in the derivation of equation 2.4, we utilized three common inverse Laplace transforms which are listed below [10, 28], and going between the third and fourth lines we made the substitution $\mu-V(\mathbf{R})=\frac{\hbar^{2}}{2m}k_{F}(\mathbf{R})^{2}$, suppressed the $\mathbf{R}$ dependent notation of $k_{F}$ so that $k_{F}=k_{F}(\mathbf{R})$, and defined $z=k_{F}s$.
@@ B.3 Gradient identities
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ ${\approx}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}.$ (2.10)
As is the case throughout this paper, an equal ($=$) sign denotes an exact identity, and an approximation sign ($\approx$) denotes an approximate one where only terms up to and including second order in derivatives of the potential $V$ (equivalently $\hbar$) are retained. Equations 2.5, 2.6 and 2.7 are straightforward to show via equation 1.3 and the definition of $z$, but equations 2.8, 2.9 and 2.10 are more complicated and deserve their own proofs. To prove equation 2.8 we use equation 2.5 and the Taylor expansion up to second order in relative coordinates $\mathbf{s}$ of $V(\mathbf{r})=V(\mathbf{R}+\mathbf{s}/2)$, given by equation 2.13. We then have
$\nabla_{\mathbf{r}}V(\mathbf{r})=\nabla_{\mathbf{r}}V(\mathbf{R}+\mathbf{s}/2){=}$ ${\frac{1}{2}\nabla_{\mathbf{R}}V(\mathbf{R}+\mathbf{s}/2)+\nabla_{\mathbf{s}}V(\mathbf{R}+\mathbf{s}/2)}$
$=$ $\frac{1}{2}\nabla_{\mathbf{R}}\bigg{(}V(\mathbf{r})+\frac{\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}}{2}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}$
$+\nabla_{\mathbf{s}}\bigg{(}V(\mathbf{r})+\frac{\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}}{2}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}.$
Taking the above expression and only keeping terms up to second derivatives of the potential $V$, we obtain
$\nabla_{\mathbf{r}}V(\mathbf{r})\approx$ $\frac{1}{2}\nabla_{\mathbf{R}}\bigg{(}V(\mathbf{r})+\frac{\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}}{2}\bigg{)}+\nabla_{\mathbf{s}}\bigg{(}\frac{\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}}{2}+\frac{1}{8}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}\bigg{)}$
$=$ $\frac{1}{2}\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$+\frac{1}{2}\nabla_{\mathbf{s}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{8}\nabla_{\mathbf{s}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}$
$=$ $\frac{1}{2}\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}+\frac{1}{2}\nabla_{\mathbf{R}}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
[…]
Next, to derive equation 2.9 we start from
@@
[…]
where we have used equation 2.6. Applying the expansion of equation 2.13 in a similar way as was done in the derivation of equation 2.8, we neglect terms with higher derivatives than second order of the potential $V$ and reduce the above expression to
$\nabla_{\mathbf{r}}^{2}V(\mathbf{r})$ $\approx\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{8}\nabla_{\mathbf{s}}^{2}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}\Big{]}+\frac{1}{2}\big{(}\nabla_{\mathbf{R}}\cdot\nabla_{\mathbf{s}}\big{)}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}$
$={\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{4}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})+\frac{1}{2}\nabla_{\mathbf{R}}^{2}V(\mathbf{R})}$
$=\nabla_{\mathbf{R}}^{2}V(\mathbf{R}).$
To derive equation 2.10 we make use of equations 2.5 and 2.8, and again only retain terms up to second order in $\nabla_{\mathbf{R}}$ to find
$\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}{\approx}$ ${\bigg{(}\frac{1}{2}\nabla_{\mathbf{R}}+\nabla_{\mathbf{s}}\bigg{)}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}+\frac{1}{2}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}}$
[…]
${+\frac{1}{2}\nabla_{\mathbf{s}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\frac{\mathbf{s}}{s}\Big{]}\cdot\frac{\mathbf{s}}{s},}$ (2.11)
To further simplify equation 2.11 into equation 2.10, we show that the second term vanishes identically by writing this term as
$\nabla_{\mathbf{s}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}$ $=\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{s_{i}}{s}\frac{\partial}{\partial s_{i}}\bigg{(}\frac{s_{j}}{s}\frac{\partial V(\mathbf{R})}{\partial R_{j}}\bigg{)}=\sum_{i=1}^{d}\sum_{j=1}^{d}\delta_{ij}\frac{s_{i}}{s^{2}}\frac{\partial V(\mathbf{R})}{\partial R_{j}}-\frac{s_{i}^{2}s_{j}}{s^{4}}\frac{\partial V(\mathbf{R})}{\partial R_{j}}$
$=\sum_{i=1}^{d}\frac{s_{i}}{s^{2}}\frac{\partial V(\mathbf{R})}{\partial R_{i}}-\sum_{j=1}^{d}\frac{s_{j}}{s^{2}}\frac{\partial V(\mathbf{R})}{\partial R_{j}}=0,$ (2.12)
while the third term of equation 2.11 simplifies to
$\frac{1}{2}\nabla_{\mathbf{s}}\Big{[}\nabla_{\mathbf{R}}\big{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\mathbf{s}\big{)}\cdot\frac{\mathbf{s}}{s}\Big{]}\cdot\frac{\mathbf{s}}{s}$ $=\sum_{i=1}^{d}\sum_{j=1}^{d}\sum_{l=1}^{d}\frac{s_{i}}{2s}\frac{\partial}{\partial s_{i}}\bigg{(}\frac{s_{j}s_{l}}{s}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{l}}\bigg{)}$
$=\sum_{i=1}^{d}\sum_{j=1}^{d}\sum_{l=1}^{d}\delta_{ij}\frac{s_{i}s_{l}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{l}}$
$\quad+\delta_{il}\frac{s_{i}s_{j}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{l}}-\frac{s_{i}^{2}s_{j}s_{l}}{2s^{4}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{l}}$
$=\sum_{i=1}^{d}\sum_{l=1}^{d}\frac{s_{i}s_{l}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{l}}$
$\quad+\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{s_{i}s_{j}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{i}}-\sum_{j=1}^{d}\sum_{l=1}^{d}\frac{s_{j}s_{l}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{j}\partial R_{l}}$
$=\sum_{i=1}^{d}\sum_{l=1}^{d}\frac{s_{i}s_{l}}{2s^{2}}\frac{\partial^{2}V(\mathbf{R})}{\partial R_{i}\partial R_{l}}=\frac{1}{2}\nabla_{\mathbf{R}}\bigg{(}\nabla_{\mathbf{R}}V(\mathbf{R})\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}.$ (2.13)
Combining equations 2.11, 2.12 and 2.13 finally yields equation 2.10.
Appendix C Alternate derivation of WKODM, KODM, and GVODM equivalence
C.1 Symmetric Bloch matrix
In references [23, 24, 25], the manifestly Hermitian and idempotent GVODM in symmetric coordinates is derived for $d=1,2,3,4$-dimensions despite the fact that it is through means of a non-Hermitian and non-idempotent WKODM method in terms of non-symmetric coordinates. In Section 2 we also proved the equivalence between the WKODM, KODM, and GVODM by applying the change of coordinates $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ to the non-symmetric Wigner-Kirkwood expansion of the Bloch density matrix. For completeness, here we will provide an additional yet equivalent $d$-dimensional derivation of the symmetric ODM equivalence which instead starts by symmetrizizing the definition of the Bloch density matrix, and then by performing the Wigner-Kirkwood expansion, now in symmetric coordinates. Starting from the same point that reference [10] does for their non-symmetric WKODM derivation, see their section 4.3, the non-symmetric Bloch density matrix is given by
Note that one can easily derive the presence of the degeneracy factor $g$ by accounting for fermionic spin within the position states and simplifying. $\hat{H}$ is the quantum one-body Hamiltonian, and we also note that the WKODM expansion has yet to be applied. Before changing to the symmetric coordinates, consider the unitary translation operator $\hat{U}(\mathbf{a})={\textrm{e}^{-(i/\hbar)\hat{\mathbf{P}}\cdot\mathbf{a}}}$, where $\hat{\mathbf{P}}=-i\hbar\nabla_{\mathbf{R}}$ is the $\mathbf{R}$-space representation of the momentum operator $\hat{\mathbf{P}}|\mathbf{p}\rangle=\mathbf{p}|\mathbf{p}\rangle$, and $\mathbf{a}$ is an arbitrary position space vector. $\hat{U}(\mathbf{a})$ acts on position space states and scalar functions $f$ as follows [30]
$\hat{U}(\mathbf{a})|\mathbf{R}\rangle$ $=|\mathbf{R}-\mathbf{a}\rangle,$
where in between the fourth and last lines we made the substitution $\hat{H}=\hat{\mathbf{P}}^{2}/(2m)+V(\mathbf{R})$. Making use of the momentum-space completeness relationship $\mathds{1}=\int d^{d}p\;|\mathbf{p}\rangle\langle\mathbf{p}|$ and the inner product relationship $\langle\mathbf{R}|\mathbf{p}\rangle=(2\pi\hbar)^{-d/2}{\textrm{e}^{(i/\hbar)\mathbf{p}\cdot\mathbf{R}}}$ allows us to recast equation 3.2 as
At this point, we can perform our symmetric semiclassical WKODM method following the same techniques used for the usual non-symmetric WKODM method within reference [10]. The process starts by making the substitution $\langle\mathbf{p}|{\textrm{e}^{-\beta\frac{\hat{\mathbf{P}}^{2}}{2m}-\beta V(\mathbf{R}+\mathbf{s}/2)}}|\mathbf{R}\rangle={\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,\mathbf{p})}}\langle\mathbf{p}|\mathbf{R}\rangle\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, where we recall from before that $H_{cl}(\mathbf{r},\mathbf{p})=\frac{\mathbf{p}^{2}}{2m}+V(\mathbf{r})$. This yields
where $\tilde{w}$ is a function which obeys
$\textrm{e}^{-\beta H_{cl}(\mathbf{R}+\mathbf{s}/2,{\mathbf{p}})-(i/\hbar)\mathbf{p}\cdot\mathbf{R}}\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)=\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta),$ (3.5)
and $\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ is a function that obeys the Bloch equation ( as defined by equation (4.35) in reference  [10])
$\frac{\partial\tilde{u}}{\partial\beta}+\hat{H}\tilde{u}=0.$ (3.6)
Here the Hamiltonian takes its position space representation, $\hat{H}=-\hbar^{2}\nabla_{\mathbf{R}}^{2}/(2m)+V(\mathbf{R})$, while $\tilde{u}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ treats both the momentum $\mathbf{p}$ and position coordinates $\mathbf{R}$ and $\mathbf{s}$ on the same footing as classical variables (opposed to operators). For simplicity, we sometimes suppress the $(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$ dependence of both the $\tilde{w}$ and $\tilde{u}$ functions, in a similar form to equation 3.6.
C.2 Symmetric WKODM expansion
Substituting equation 3.5 into equation 3.6, making the second order semiclassical approximation $\tilde{w}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)\approx 1+\hbar\tilde{w}_{1}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)+\hbar^{2}\tilde{w}_{2}(\mathbf{R},\mathbf{s},\mathbf{p};\beta)$, and collecting terms based on their order of $\hbar$ while dividing through by a common factor of $\textrm{e}^{-\beta H_{\text{cl}}(R+s/2,p)-(i/\hbar)p\cdot R}$ yields